OpenStudy (anonymous):

Find the vector, given its magnitude and direction angle. absolute value of u= 4 and theta= 119 degrees

OpenStudy (aum):

\(u\cos(\theta)i + u\sin(\theta)j\)

OpenStudy (anonymous):

im confused what does tht mean?

OpenStudy (aum):

Substitute u = 4 and \(\theta = 119\) in the above equation and simplify.

OpenStudy (anonymous):

oh ok thanks

OpenStudy (aum):

The x-component of the vector is \((u\cos\theta)\). The y-component of the vector is \((u\sin\theta)\).

OpenStudy (anonymous):

kk thaniks